MrQuestion:

The value of 23 + 42 = ___.

supie:

You can use the calculator feature below.

MrQuestion:

Its supposed to say

MrQuestion:

2 to the power of 3

supie:

oh

MrQuestion:

And 4 to the power of 2

snowflake0531:

2^3 + 4^2 2 to the power of 3 is 8, 4 the power of 3 is 16, add em

supie:

\(2^3 + 4^2\) like that?

MrQuestion:

Ye

supie:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @supie \(2^3 + 4^2\) like that? \(\color{#0cbb34}{\text{End of Quote}}\) \(2^3 + 4^2=2 \times 2\times 2 +4 \times 4 \)

MrQuestion:

I was cofused because my teacher makes everything complicated

MrQuestion:

thx

supie:

np

supie:

i dont think that's the answer tho- just incase you thought it was the answer ._.

AZ:

So what you need to know about exponents is- \(a^b\) means you're multiplying a by itself b times so like 4^2 or \( 4^2\) means 4 times 4 4^5 or \( 4^5\) means 4 * 4 * 4 * 4 * 4 so four multiplied by itself 5 times